OpenStudy (anonymous):

Factor the trinomial: 6x2 + 7x - 20 A. (3x + 6)(x + 1) B. (3x + 2)(2x - 10) C. (2x - 5)(3x + 4) D. (3x - 4)(2x + 5)

OpenStudy (hoblos):

6x2 + 7x - 20 = 6x2 +15x - 8x -20 = 3x(2x+5) -4(2x+5) = (2x+5)(3x-4)